# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v1

@@ A.2 Zeroth order Self-energy
$\Phi[G]$ $=\frac{1}{2}G_{o_{3}^{\prime},o_{1}^{\prime}}(r_{3}^{\prime},r_{1}^{\prime})\delta_{\sigma_{1}^{\prime}s_{3}^{\prime}}(\Gamma_{o_{1}^{\prime}o_{2}^{\prime}o_{3}^{\prime}o_{4}^{\prime}}\delta_{\sigma_{1}^{\prime}s_{3}^{\prime}}\delta_{\sigma_{2}^{\prime}s_{4}^{\prime}}-\Gamma_{o_{1}^{\prime}o_{2}^{\prime}o_{4}^{\prime}o_{3}^{\prime}}\delta_{\sigma_{1}^{\prime}s_{4}^{\prime}}\delta_{\sigma_{2}^{\prime}s_{3}^{\prime}})\delta_{(r_{1}^{\prime},r_{2}^{\prime},r_{3}^{\prime},r_{4}^{\prime})}G_{o_{4}^{\prime},o_{2}^{\prime}}(-r_{4}^{\prime},r_{2}^{\prime})\delta_{\sigma_{2}^{\prime}s_{4}^{\prime}}$ (93)
$=\frac{1}{2}G_{o_{3}^{\prime},o_{1}^{\prime}}(r_{1}^{\prime},r_{1}^{\prime})(2\Gamma_{o_{1}^{\prime}o_{2}^{\prime}o_{3}^{\prime}o_{4}^{\prime}}-\Gamma_{o_{1}^{\prime}o_{2}^{\prime}o_{4}^{\prime}o_{3}^{\prime}})G_{o_{4}^{\prime},o_{2}^{\prime}}(r_{1}^{\prime},r_{1}^{\prime})\,.$ (94)
@@
As discussed elsewhere [78, 88] such a form of the Luttinger-Ward functional induces a local and static self-energy, which has to be accounted for in our calculations. We recall the definition of the self energy
$\Sigma_{o_{1},o_{3}}(r_{1},r_{3})=\frac{\delta\Phi[G]}{\delta G_{o_{3},o_{1}}(r_{3},r_{1})}=(2\Gamma_{o_{1}^{\prime},o_{2}^{\prime},o_{3}^{\prime},o_{4}^{\prime}}-\Gamma_{o_{1}^{\prime},o_{2}^{\prime},o_{4}^{\prime},o_{3}^{\prime}})G_{o_{4}^{\prime},o_{2}^{\prime}}\delta_{r_{1},r_{3}}\,,$ (95)

### v3

@@ A.2 Zeroth order Self-energy
$\Phi[G]$ $=\frac{1}{2}G_{o_{3^{\prime}},o_{1^{\prime}}}(\boldsymbol{\tau}_{1}^{\prime},\boldsymbol{\tau}_{1}^{\prime})(2\Gamma_{o_{1^{\prime}}o_{2^{\prime}}o_{3^{\prime}}o_{4^{\prime}}}-\Gamma_{o_{1^{\prime}}o_{2^{\prime}}o_{4^{\prime}}o_{3^{\prime}}})G_{o_{4^{\prime}},o_{2^{\prime}}}(\boldsymbol{\tau}_{1}^{\prime},\boldsymbol{\tau}_{1}^{\prime})\,.$ (98)
@@
For a diagrammatic representation of the Luttinger-Ward functional within the TPSC approximation we refer the reader to Ref. [80] as such a representation is identical between single and multi-orbital. As discussed elsewhere [80, 90, 91] such a form of the Luttinger-Ward functional induces a local and static self-energy, which has to be accounted for in our calculations. We recall the definition of the self energy
$\Sigma_{o_{1},o_{3}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3})=\frac{\delta\Phi[G]}{\delta G_{o_{3},o_{1}}(\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{1})}=(2\Gamma_{o_{1},o_{2^{\prime}},o_{3},o_{4^{\prime}}}-\Gamma_{o_{1},o_{2^{\prime}},o_{4^{\prime}},o_{3}})G_{o_{4^{\prime}},o_{2^{\prime}}}(\boldsymbol{\tau}_{1}^{\prime},\boldsymbol{\tau}_{1}^{\prime})\delta_{\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3}}\,.$ (99)